OpenStudy (anonymous):

What is the value of x rounded to the nearest hundredth? Type the numeric answer only in the box below.

OpenStudy (anonymous):

@waterineyes

OpenStudy (anonymous):

Just do the same as we did earlier..

OpenStudy (anonymous):

@waterineyes how would i set it up?

OpenStudy (anonymous):

\[\sin(49) = \frac{Perpendicular}{Hypotenuse}\] Try to put in the values given to you.. :)

OpenStudy (anonymous):

i got 0.0137;/

OpenStudy (anonymous):

thats not right... i dont understand

OpenStudy (anonymous):

Tell me what is here perpendicular's length??

OpenStudy (anonymous):

55

OpenStudy (anonymous):

So, your concepts about which is perpendicular, base and hypotenuse are weak..:)

OpenStudy (anonymous):

Give me time to explain that..

OpenStudy (anonymous):

would it be 41.51... I dont know. I think this is wrong... or 41.50

OpenStudy (anonymous):

In a Right Angled Triangle, The side opposite to angle is perpendicular.. 

OpenStudy (anonymous):



OpenStudy (anonymous):

The side opposite to angle 49 is your perpendicular which is x.. Hypotenuse is 55. \[\sin(49) = \frac{x}{55} \implies x = \sin(49) \times 55\]

OpenStudy (anonymous):

\(x = 41.5\) is right..

OpenStudy (anonymous):

and is that rounded to the nearest hundredth @waterineyes

OpenStudy (anonymous):

No..

OpenStudy (anonymous):

so what would it be;/ because i placed 41.51 and it was wrong;/

OpenStudy (anonymous):

\(x = 41. 51\)

OpenStudy (anonymous):

How is it wrong??

OpenStudy (anonymous):

i have no idea...

OpenStudy (anonymous):

Nearest hundred of .509 would be??

OpenStudy (anonymous):

I have no Idea too.. :)